# and quantum information science

### v2

@@ 7.6 Nielsen’s complexity and holography
There are also attempts to understand holographic complexity by exploiting equivalence between bulk and boundary symplectic forms [388, 48, 30], inspired by behavior under conformal transformations in Bañados geometries [136, 389]. [^38] Alternative attempts, where gravity arises as a consequence of spacetime optimizing the computational cost of its own dynamics, are discussed in [390, 41, 391].
The relation to symplectic forms also motivated the connection between quantum complexity in CFTs and gravity for the Fubini-Study cost function obtained in section 5.6.3. To understand this relation, we observe that one can map trajectories built from the unitary operators (56) of a CFTd to the geodesics of a massive particle of mass $m$ in AdSd+1 spacetime. The Fubini-Study metric (57b) receives a natural geometric interpretation in terms of the minimal $\delta X_{\rm min}$ and maximal $\delta X_{\rm max}$ perpendicular distances between infinitesimally nearby geodesics (as illustrated in Fig. 1 of Ref. [42])
@@
It would be interesting to generalize the relation (141) to complexity associated with states that are related by a finite transformation. Furthermore, the connection to cost functions other than the Fubini-Study one is a subject of an ongoing investigation [140].
For general excited states in holography, we must deal with two distinct bulk geometries whose direct comparison should occur via covariant notions (for example, through one of the holographic complexity proposals) associated with each of the corresponding asymptotic boundaries. The authors of  [44] proposed the following idea. For circuits in which the circuit parameter is the physical time in QFT, one layer of the circuit is simply given by the instantaneous QFT Hamiltonian. This Hamiltonian is defined in terms of the energy-momentum tensor components smeared along a time slice, e.g. $H=\int d^{d-1}x\,T_{tt}(\vec{x})$. The infinitesimal increment of the Fubini-Study distance (57b) is proportional to the variance of the Hamiltonian in an instantaneous state. However, given that the Hamiltonian is a sum of local operators, its variance is given by a sum of two-point functions of local operators. In AdS/CFT, the holographic dictionary [365, 366, 392, 393] provides a systematic geometric way of computing correlation functions of the boundary QFTs. As a result, the Fubini-Study cost (57b) associated with one layer of time evolution on the boundary is always geometric by the virtue that all its ingredients are geometric. This statement therefore expresses an exact mapping between two geometries: the auxiliary complexity geometry of the Fubini-Study cost on one hand, and the gravitational geometry of the bulk on the other. While the above statement is geometric, the bulk manifestation of the Fubini-Study cost that leads to it is rather complicated and implicit. However, in the context of holographic two-dimensional CFTs and states represented holographically as solutions of the Einstein’s equations with negative cosmological constant in three bulk dimensions, one can relate two-point functions of the energy-momentum tensor to geodesic lengths in the bulk [44]. This leads to a much more explicit expression for the Fubini-Study cost and a direct connection to the kinematic space program [394, 395, 396].
@@ 7.7 Computational pseudorandomness and the holographic dictionary
In all the above discussions, we mainly considered the volume of wormholes as a dual description of some notion of the complexity of time-evolving thermofield double state. It was observed by Susskind [247] that the steady growth of a wormhole’s volume poses a challenge in the context of the AdS/CFT correspondence: The dual quantity to the wormhole’s volume cannot be a local observable as those are expected to equilibrate and hence saturate quickly in a quantum theory. Instead, [247] proposed that the dual of the wormhole’s volume is a global property of the dual state such as quantum complexity. However, from a complexity-theoretic point of view, there appears to be a mismatch between these two quantities. The volume of a wormhole is a simple quantity that can be easily calculated from the gravitational description. In comparison, quantum complexity is notoriously difficult to compute or even bound. Making this mismatch concrete, Ref. [397] argues that either quantum gravity cannot be simulated efficiently on a quantum computer, or some entries in the holographic dictionary require exponentially long computations. The former would violate the extended Church-Turing thesis – the widely held belief that all physical processes can be efficiently simulated on a quantum computer. More precisely, Ref. [397] models the dual CFT as a quantum system initialized in a thermofield-double state $|\mathrm{TFD}\rangle$ dual to a zero-volume wormhole. This state is undergoing time evolution under a fixed Hamiltonian $H_{\mathrm{CFT}}$ such that $|\mathrm{TFD}\rangle$ is a low-energy state for $H_{\mathrm{CFT}}$. However, in this time evolution, the system undergoes a number of random ‘shocks’ acting locally on single qubits:
@@
In Ref. [397], the authors conjecture that the time-evolution with shocks in Eq. (142), while possible to efficiently simulate on a quantum computer, is computationally indistinguishable from a Haar random state. Such an asymmetry between the easiness of preparing a state and the hardness of distinguishing it from a random state is called quantum pseudorandomness [398, 399]. More precisely, an ensemble of states, parameterized by a random key $k$ is called pseudorandom if 1) it can be efficiently prepared by polynomial-sized quantum circuits and 2) any quantum computer requires superpolynomial time to distinguish the ensemble from Haar random states. This implies that even if the CFT evolution under shocks is computationally simple, its output states may appear indistinguishable from truly random states, making certain bulk properties computationally inaccessible from the CFT side. In particular, quantum circuit complexity cannot be computable using only polynomial-time quantum computations and measurements, as such computations would be able to rapidly distinguish pseudorandom and random states, which should be impossible. The existence of pseudorandom states was proven in Ref. [398] under mild cryptographic assumptions.
@@
We can now consider two time evolutions as in Eq. (142) for vastly different times, both undergoing shocks. Pseudorandomness implies that the two resulting states cannot be efficiently distinguished from Haar random states and, therefore, from each other. However, the two states will correspond to wormholes of different volumes in any holographic dictionary, and the volume is efficiently computable from the AdS state using only coarse-grained properties of the metric. Ref. [397] argues that this tension can only be resolved if either 1) quantum gravity cannot be efficiently simulated on a quantum computerm or 2) the holographic dictionary cannot be computed in polynomial time. Notice that this argument actually does not require the Volume$=$Complexity conjecture. In fact, the existence of pseudorandomness implies that any quantity that distinguishes between long and short time evolutions under shocks must be computationally hard to estimate.
@@
A central concept in the AdS/CFT correspondence is the relation between entanglement entropy of subregions in the CFT and the area of a corresponding extremal surface in AdS space, as specified by the Ryu-Takayanagi formula [5]. The existence of this formula suggests that there may be a way of reconstructing AdS geometry from patterns of entanglement in the CFT – a task that we may call geometric reconstruction (see, e.g., [400, 401, 402, 403] for earlier approach to reconstructing local operators). With this in mind, suppose that we consider CFT states that satisfy the restrictive holographic entropy cone inequalities [404]. These inequalities define a cone in the space spanned by the entanglement entropies of subregions and are a necessary requirement for geometrization of entanglement by the Ryu-Takayanagi formula. It was shown in Ref. [405] that “geometric reconstruction” is hard even for states whose entanglement entropy satisfies the holographic entropy cone. Similar to the argument made in Ref. [397] this hardness result uses pseudorandomness and, more precisely, the concept of pseudoentanglement. Pseudoentanglement [406, 407] refers to the phenomenon that even exponentially large gaps in entanglement can be computationally hard to detect [406], which indicates that entanglement is an “unfeelable” quantity, similar to quantum circuit complexity. In particular, Ref. [405] uses recently constructed low-entangling pseudorandom unitaries [408] to show that two states with arbitrarily different bulk geometries can correspond to indistinguishable states in the CFT. As the construction in Ref. [405] does not require event horizons, the result suggests that reconstruction of the bulk geometry might even be hard outside the event horizon. Previously, the presence of an event horizon was identified [409] as a way to reconcile the Church-Turing thesis with the duality of volume and complexity: an observer that needs to cross the event horizon to estimate the volume will then be unable to communicate the findings to an outside observer. This would be consistent with an exponentially hard-to-estimate dual quantity, such as complexity. Likewise, the authors of [236] argued that when the time evolution operator of a holographic CFT is approximately pseudorandom (or Haar random) on some code subspace, then there must be an event horizon in the corresponding bulk dual. The apparent tension with [405] may be resolved by the fact that the construction in the latter does not take the CFT Hamiltonian into account and produces high-energy states.
@@
These results raise questions about the broader landscape of computational complexity in holography, particularly in relation to the Python’s Lunch conjecture [410]. This conjecture posits that if a geometry contains a local minimal surface that is not globally minimal, then finding the CFT operator corresponding to a local bulk operator inserted in the region surrounded by the local minimal surface is exponentially hard. The conjecture suggests that if the bulk contains locally minimal but not globally minimal surfaces, then efficient application of operators in the CFT is exponentially complex. The strong Python’s lunch conjecture then further posits that locally minimal surfaces are the only obstruction to efficient “operator reconstruction” [405]. A follow-up [231] shows that the strong Python’s lunch conjecture is compatible with the findings of Ref. [397] on the reconstruction of a wormhole’s volume. Similarly, [411, 412] show that it is exponentially hard to decode the state of a field in the interior of a black hole from the distant Hawking radiation entangled with it, as expected from the Hayden-Preskill scenario [413]. The results of Ref. [405] then suggest that the complexity of operator reconstruction in the CFT does not always align with the complexity of bulk geometry reconstruction. These questions about the complexity of the holographic dictionary, inspired by the volume=complexity, require and deserve future investigations.
@@
Finally, recent developments relate the growth of the volume of Einstein-Rosen bridges to the randomness of the underlying microscopic quantum gravity state [90, 91]. Concretely, the authors of [90, 91] found gravitational duals to continuous random circuits (see Sec. 4.5 for a survey of random circuits in the context of circuit complexity). The resulting geometries are long wormholes with matter inhomogeneities, dubbed ER-caterpillars. The construction allowed the authors to demonstrate a relation between the length of the ER bridge and the degree of microscopic randomness of the state as a function of time. More precisely, [90, 91] showed that an ER-caterpillar of length $l$ has an underlying quantum state well-approximated by a quantum state k-design with $k\propto l$ when $l$ is large. It would be interesting to understand the implications of this construction in the context of the pseudorandomness mentioned above.
@@ 8 Paradigms for complexity III: tensor-network-inspired definitions of complexity
Path-integral complexity [38, 39] was the first approach to measuring complexity in quantum field theories that was motivated by holography. This approach arose in the context of tensor networks, which provide variational ansatze for many-body quantum wavefunctions. In particular, optimized tensor networks such as, e.g., MERA [17, 414] or its continuous counterpart cMERA [18, 98], give rise to a geometric structure that reflects the entanglement patterns of a quantum state. An analogy between the tensor network approach to constructing quantum states and the structure of the holographic duality had already been noted by Swingle [239]. This proposal seeded a new line of research that led Susskind [6] to argue that the growth of the volume of Einstein-Rosen bridges should be related to complexity, estimated by the number of tensors (the effective volume of the network) in the tensor-network representation of the time-evolved thermofield-double state [245]. In Sec. 8.1, we review a notion of complexity in CFT called path-integral complexity inspired by the parallelism with tensor networks. In Sec. 8.2 we discuss the holographic interpretation of this quantity in AdS spacetime.
@@ 8.1 Complexity from path-integral optimization
Motivated by these developments, Ref. [38] aimed at extending intuitions from the study of tensor networks to general, strongly-interacting QFTs and to use these ideas to define a natural measure of complexity using path integrals. A brief review of their construction follows.
@@
The main new idea in [38] was to introduce a non-trivial background metric on the space where the path integral (143) is performed (keeping the boundary conditions for the fields fixed), and to interpret this metric as an (unoptimized) density of gates of a continuous tensor network. For our 2D CFT example, all the metrics can be put into the Weyl-flat form parametrized by a scalar function $\phi(\tau,x)$ as follows:
@@
Figure 13: A cartoon explaining path-integral optimization in CFTs as an analog of tensor-network optimization. Figure from [38] with Euclidean time denoted as $z=-\tau$.
@@
[…]
@@
with the boundary condition (144). Recall that we are interpreting the metric as encoding the density of gates in the tensor network. This procedure is therefore called the path-integral optimization.
@@
Similar optimization was also defined for the vacuum and thermal states of higher-dimensional CFTs [39], but with a prescription to optimize only over the Weyl factor of the background metric on which the CFT path integral is evaluated (unlike in 2D, this is only a subset of possible geometries in higher-dimensions). Again, it was found that optimal metrics have a constant negative Ricci scalar curvature i.e., are hyperbolic.
@@
Another idea from [38, 39], dubbed path-integral complexity, was to interpret the Liouville action as a novel measure of complexity for wave functions prepared using Euclidean path integrals. The main reason to advocate this interpretation was that the on-shell Liouville action is proportional to the volume of the space on which the path integral is computed. Indeed, after fixing the normalization such that the integral of the Weyl factor $e^{2\phi}$ over a unit area (a square region of size $\epsilon$ in the original metric) has a value 1 (see detailed discussion in [415]), we can interpret the volume of the Euclidean space as the number of tensors in the optimal network. The on-shell action associated with the CFT vacuum state reads
[…]
@@
This result is also consistent with qualitative expectations and holographic conjecture [25], which estimates the leading divergence in complexity of the CFT state by the spatial volume.
@@
Later on, researchers extended the path-integral optimization method to other settings. As we already mentioned, [38, 39] proposed a higher-dimensional generalization of the action (148). The the optimization of 2D CFTs deformed by relevant operators was studied in [415], where a modified Liouville action was minimized over both metrics and local deformation couplings. The authors of [416] investigated path integrals for inhomogeneous CFTs, that are usually interpreted as 2D CFTs on non-trivial background metrics. From the perspective of the optimization, the starting point is the CFT on a non-trivial background metric ($\hat{g}_{ab}$ instead of $\delta_{ab}$) and then consider $g_{ab}=e^{2\phi}\hat{g}_{ab}$ as the metric for optimization. This leads to additional background Ricci curvature term $\hat{R}$ in the Liouville equation which can nevertheless be solved using standard uniformization techniques.
@@
Soon after [38], reference [417] argued that a variation of path integral complexity in holographic 2D CFTs leads to Einstein’s equations in the bulk of AdS3. Moreover, the Liouville action (or its covariant version, called the Polyakov action) motivated the treatment of Nielsen’s complexity by introducing trajectories on the unitary manifold generated by sequences of 2D conformal transformations [40] (see Sec. 5.6.2).
@@
The formulation in terms of a Liouville action also inspired a series of developments in the tensor-network community. They used the path integral approach to cMERA to define tensor networks in terms of the background geometry on which the state is prepared [418, 419]. In particular, they showed how the background metric parameters translate into details of the tensor network. This idea was later used in [420] to argue for a Nielsen-type interpretation of the Liouville action in terms of an appropriate choice of the cost function on the complexity geometry. As opposed to the developments on Nielsen’s complexity discussed in Sec. 5, the work [420] necessarily utilized circuits that contained both unitary and Hermitian gates associated with components of the energy-momentum tensor operator. This is similar in spirit to the results discussed in Sec. 7.5.1 of [370] and has the same origin: Euclidean time evolution is an efficient and natural method for preparing useful states. This points towards a mathematical generalization of the Nielsen’s complexity that allows for exponentiation of anti-Hermitian (giving rise to unitary transformations), as well as Hermitian matrices (giving rise in particular to Euclidean time evolution).
@@
Finally, the results in Ref. [420] pointed out that a general cost function for path integral optimization will also contain positive powers of the cut-off $\epsilon$ associated with higher derivatives of $\phi$ in the spirit of an effective field theory. However, the solutions of Eq. (147) acquire derivatives that are themselves of the order of $1/\epsilon$, breaking the hierarchy within these more general cost functions that left the contribution (146) as the dominant term. Thus we need a more systematic inclusion of the UV cut-off $\epsilon$, bringing us to the subject of the next section.

### v3

@@ 7.6 Nielsen’s complexity and holography
There are also attempts to understand holographic complexity by exploiting equivalence between bulk and boundary symplectic forms Belin:2018fxe, Belin:2018bpg, Belin:2022xmt, inspired by behavior under conformal transformations in Bañados geometries Flory:2018akz, Flory:2019kah. [^41] Alternative attempts, where gravity arises as a consequence of spacetime optimizing the computational cost of its own dynamics, are discussed in Pedraza:2022dqi, Chandra:2021kdv, Carrasco:2023fcj.
The relation to symplectic forms also motivated the connection between quantum complexity in CFTs and gravity for the Fubini-Study cost function obtained in section 5.6.3. To understand this relation, we observe that one can map trajectories built from the unitary operators (56) of a CFTd to the geodesics of a massive particle of mass $m$ in AdSd+1 spacetime. The Fubini-Study metric (57b) receives a natural geometric interpretation in terms of the minimal $\delta X_{\rm min}$ and maximal $\delta X_{\rm max}$ perpendicular distances between infinitesimally nearby geodesics (as illustrated in Fig. 1 of Ref. Chagnet:2021uvi)
@@
It would be interesting to generalize the relation (141) to complexity associated with states that are related by a finite transformation. Furthermore, the connection to cost functions other than the Fubini-Study one is a subject of an ongoing investigation CFTpen.
The previous approach described the entire circuit (starting from a scalar primary state and employing CFT generators as gates) in terms of a massive particle trajectory in AdS. This approach does not immediately lend itself to an intuitive AdS/CFT interpretation in which quantum states are viewed as living on the boundary of AdS space. For this reason, it is desirable to seek a formulation of the state–complexity problem for general excited states in holography, where the states along the circuit can be more directly mapped to the boundary of AdS. Naively, the holographic complexity picture would require two distinct bulk geometries for the reference and target states (and, in fact, for any state along the circuit), since gravity would dictate the state’s time evolution in a way not necessarily aligned with its evolution along the circuit. However, it turns out that in certain cases it is possible to embed both the reference and target states in the same holographic geometry and connect them through unitary time evolution on the boundary, which can be viewed as a circuit in physical time. Although finding such a geometry is not straightforward in general, the problem simplifies when considering two states that are very close to each other. This is precisely the setting of Erdmenger:2022lov, which proposed the following idea.
For circuits in which the circuit parameter is the physical time in QFT, one layer of the circuit is simply given by the instantaneous QFT Hamiltonian. This Hamiltonian is defined in terms of the energy-momentum tensor components smeared along a time slice, e.g. $H=\int d^{d-1}x\,T_{tt}(\vec{x})$. The infinitesimal increment of the Fubini-Study distance (57b) is proportional to the variance of the Hamiltonian in an instantaneous state. However, given that the Hamiltonian is a sum of local operators, its variance is given by a sum of two-point functions of local operators. In AdS/CFT, the holographic dictionary Gubser:1998bc, Witten:1998qj, Skenderis:2008dh, Skenderis:2008dg provides a systematic geometric way of computing correlation functions of the boundary QFTs. As a result, the Fubini-Study cost (57b) associated with one layer of time evolution on the boundary is always geometric by the virtue that all its ingredients are geometric. This statement therefore expresses an exact mapping between two geometries: the auxiliary complexity geometry of the Fubini-Study cost on one hand, and the gravitational geometry of the bulk on the other. While the above statement is geometric, the bulk manifestation of the Fubini-Study cost that leads to it is rather complicated and implicit. However, in the context of holographic two-dimensional CFTs and states represented holographically as solutions of the Einstein’s equations with negative cosmological constant in three bulk dimensions, one can relate two-point functions of the energy-momentum tensor to geodesic lengths in the bulk Erdmenger:2022lov. This leads to a much more explicit expression for the Fubini-Study cost and a direct connection to the kinematic space program Czech:2015qta, Czech:2016xec, deBoer:2016pqk.
@@ 7.7 Computational pseudorandomness and the holographic dictionary
In all the above discussions, we mainly considered the volume of wormholes as a dual description of some notion of the complexity of time-evolving thermofield double state. It was observed by Susskind Susskind:2014rva that the steady growth of a wormhole’s volume poses a challenge in the context of the AdS/CFT correspondence: the dual quantity to the wormhole’s volume cannot be a local observable as those are expected to equilibrate and hence saturate quickly in a quantum theory. Instead, Susskind:2014rva proposed that the dual of the wormhole’s volume is a global property of the dual state such as quantum complexity. However, from a complexity-theoretic point of view, there appears to be a mismatch between these two quantities. The volume of a wormhole is a simple quantity that can be easily calculated from the gravitational description. In comparison, quantum complexity is notoriously difficult to compute or even bound. Making this mismatch concrete, Ref. Bouland:2019pvu argues that either quantum gravity cannot be simulated efficiently on a quantum computer, or some entries in the holographic dictionary require exponentially long computations. The former would violate the extended Church-Turing thesis – the widely held belief that all physical processes can be efficiently simulated on a quantum computer. More precisely, Ref. Bouland:2019pvu models the dual CFT as a quantum system initialized in a thermofield-double state $|\mathrm{TFD}\rangle$ dual to a zero-volume wormhole. This state is undergoing time evolution under a fixed Hamiltonian $H_{\mathrm{CFT}}$ such that $|\mathrm{TFD}\rangle$ is a low-energy state for $H_{\mathrm{CFT}}$. However, in this time evolution, the system undergoes a number of random ‘shocks’ acting locally on single qubits:
@@
In Ref. Bouland:2019pvu, the authors conjecture that the time-evolution with shocks in Eq. (142), while possible to efficiently simulate on a quantum computer, is computationally indistinguishable from a Haar random state. Such an asymmetry between the easiness of preparing a state and the hardness of distinguishing it from a random state is called quantum pseudorandomness ji2018pseudorandom, ma2024construct. More precisely, an ensemble of states, parameterized by a random key $k$ is called pseudorandom if 1) it can be efficiently prepared by polynomial-sized quantum circuits and 2) any quantum computer requires superpolynomial time to distinguish the ensemble from Haar random states. This implies that even if the CFT evolution under shocks is computationally simple, its output states may appear indistinguishable from truly random states, making certain bulk properties computationally inaccessible from the CFT side. In particular, quantum circuit complexity cannot be computable using only polynomial-time quantum computations and measurements, as such computations would be able to rapidly distinguish pseudorandom and random states, which should be impossible. The existence of pseudorandom states was proven in Ref. ji2018pseudorandom under mild cryptographic assumptions.
@@
We can now consider two time evolutions as in Eq. (142) for vastly different times, both undergoing shocks. Pseudorandomness implies that the two resulting states cannot be efficiently distinguished from Haar random states and, therefore, from each other. However, the two states will correspond to wormholes of different volumes in any holographic dictionary, and the volume is efficiently computable from the AdS state using only coarse-grained properties of the metric. Ref. Bouland:2019pvu argues that this tension can only be resolved if either 1) quantum gravity cannot be efficiently simulated on a quantum computer or 2) the holographic dictionary cannot be computed in polynomial time. Notice that this argument actually does not require the Volume==Complexity conjecture. In fact, the existence of pseudorandomness implies that any quantity that distinguishes between long and short time evolutions under shocks must be computationally hard to estimate.
@@
A central concept in the AdS/CFT correspondence is the relation between entanglement entropy of subregions in the CFT and the area of a corresponding extremal surface in AdS space, as specified by the Ryu-Takayanagi formula Ryu:2006bv. The existence of this formula suggests that there may be a way of reconstructing AdS geometry from patterns of entanglement in the CFT – a task that we may call geometric reconstruction (see, e.g., Hamilton:2005ju, Hamilton:2006az, Hamilton:2006fh, Kabat:2011rz for earlier approach to reconstructing local operators). With this in mind, suppose that we consider CFT states that satisfy the restrictive holographic entropy cone inequalities bao2015holographic. These inequalities define a cone in the space spanned by the entanglement entropies of subregions and are a necessary requirement for geometrization of entanglement by the Ryu-Takayanagi formula. It was shown in Ref. akers2024holographic that “geometric reconstruction” is hard even for states whose entanglement entropy satisfies the holographic entropy cone. Similar to the argument made in Ref. Bouland:2019pvu this hardness result uses pseudorandomness and, more precisely, the concept of pseudoentanglement. Pseudoentanglement aaronson2022quantum, gheorghiu2020estimating refers to the phenomenon that even exponentially large gaps in entanglement can be computationally hard to detect aaronson2022quantum, which indicates that entanglement is an “unfeelable” quantity, similar to quantum circuit complexity. In particular, Ref. akers2024holographic uses recently constructed low-entangling pseudorandom unitaries schuster2024random to show that two states with arbitrarily different bulk geometries can correspond to indistinguishable states in the CFT. As the construction in Ref. akers2024holographic does not require event horizons, the result suggests that reconstruction of the bulk geometry might even be hard outside the event horizon. Previously, the presence of an event horizon was identified susskind2020horizons as a way to reconcile the Church-Turing thesis with the duality of volume and complexity: an observer that needs to cross the event horizon to estimate the volume will then be unable to communicate the findings to an outside observer. This would be consistent with an exponentially hard-to-estimate dual quantity, such as complexity. Likewise, the authors of Engelhardt:2024hpe argued that when the time evolution operator of a holographic CFT is approximately pseudorandom (or Haar random) on some code subspace, then there must be an event horizon in the corresponding bulk dual. The apparent tension with akers2024holographic may be resolved by the fact that the construction in the latter does not take the CFT Hamiltonian into account and produces high-energy states.
@@
These results raise questions about the broader landscape of computational complexity in holography, particularly in relation to the Python’s Lunch conjecture brown2020python. This conjecture posits that if a geometry contains a local minimal surface that is not globally minimal, then finding the CFT operator corresponding to a local bulk operator inserted in the region surrounded by the local minimal surface is exponentially hard. The conjecture suggests that if the bulk contains locally minimal but not globally minimal surfaces, then efficient application of operators in the CFT is exponentially complex. The strong Python’s lunch conjecture then further posits that locally minimal surfaces are the only obstruction to efficient “operator reconstruction” akers2024holographic. A follow-up engelhardt2022finding shows that the strong Python’s lunch conjecture is compatible with the findings of Ref. Bouland:2019pvu on the reconstruction of a wormhole’s volume. Similarly, Gyongyosi:2022vaf, Balasubramanian:2022fiy show that it is exponentially hard to decode the state of a field in the interior of a black hole from the distant Hawking radiation entangled with it, as expected from the Hayden-Preskill scenario Yoshida:2017aa. The results of Ref. akers2024holographic then suggest that the complexity of operator reconstruction in the CFT does not always align with the complexity of bulk geometry reconstruction. These questions about the complexity of the holographic dictionary, inspired by the volume=complexity, require and deserve future investigations.
@@
Finally, recent developments relate the growth of the volume of Einstein-Rosen bridges to the randomness of the underlying microscopic quantum gravity state Magan:2024aet, Magan:2025hce. Concretely, the authors of Magan:2024aet, Magan:2025hce found gravitational duals to continuous random circuits (see Sec. 4.5 for a survey of random circuits in the context of circuit complexity). The resulting geometries are long wormholes with matter inhomogeneities, dubbed ER-caterpillars. The construction allowed the authors to demonstrate a relation between the length of the ER bridge and the degree of microscopic randomness of the state as a function of time. More precisely, Magan:2024aet, Magan:2025hce showed that an ER-caterpillar of length $l$ has an underlying quantum state well-approximated by a quantum state k-design with $k\propto l$ when $l$ is large. It would be interesting to understand the implications of this construction in the context of the pseudorandomness mentioned above.
@@ 8 Paradigms for complexity III: tensor-network-inspired definitions of complexity
Path-integral complexity Caputa:2017urj, Caputa:2017yrh was the first approach to measuring complexity in quantum field theories that was motivated by holography. This approach arose in the context of tensor networks, which provide variational ansatze for many-body quantum wavefunctions. In particular, optimized tensor networks such as, e.g., MERA Vidal:2007hda, Vidal:2008zz or its continuous counterpart cMERA Haegeman:2011uy, Nozaki:2012zj, give rise to a geometric structure that reflects the entanglement patterns of a quantum state. An analogy between the tensor network approach to constructing quantum states and the structure of the holographic duality had already been noted by Swingle Swingle:2009bg. This proposal seeded a new line of research that led Susskind Susskind:2014moa to argue that the growth of the volume of Einstein-Rosen bridges should be related to complexity, estimated by the number of tensors (the effective volume of the network) in the tensor-network representation of the time-evolved thermofield-double state Hartman:2013aa. In Sec. 8.1, we review a notion of complexity in CFT called path-integral complexity inspired by the parallelism with tensor networks. In Sec. 8.2 we discuss the holographic interpretation of this quantity in AdS spacetime.
@@ 8.1 Complexity from path-integral optimization
Motivated by these developments, Ref. Caputa:2017urj aimed at extending intuitions from the study of tensor networks to general, strongly-interacting QFTs and to use these ideas to define a natural measure of complexity using path integrals. A brief review of their construction follows.
@@
The main new idea in Caputa:2017urj was to introduce a non-trivial background metric on the space where the path integral (143) is performed (keeping the boundary conditions for the fields fixed), and to interpret this metric as an (unoptimized) density of gates of a continuous tensor network. For our 2D CFT example, all the metrics can be put into the Weyl-flat form parametrized by a scalar function $\phi(\tau,x)$ as follows:
@@
Figure 13: A cartoon explaining path-integral optimization in CFTs as an analog of tensor-network optimization. Figure from Caputa:2017urj with Euclidean time denoted as $z=-\tau$.
@@
$S_{\mathcal{L}}[\phi]=\frac{c}{24\pi}\int^{\infty}_{-\infty}\!dx\!\int^{0}_{-\infty}\!d\tau\!\left[(\partial_{x}\phi)^{2}+(\partial_{\tau}\phi)^{2}+\mu e^{2\phi}\right],\penalty 10000\ \penalty 10000\$(146)
@@
with the boundary condition (144). This procedure is called the path-integral optimization. Recall that we are interpreting the metric as encoding the density of gates in the tensor network. One attempt to make this precise showed a correspondence between unitary and isometric tensors of the (c)MERA network and the kinetic and potential terms in the Liouville action Caputa:2017yrh, Czech:2017ryf. Likewise, Camargo:2019isp, Caputa:2018kdj argued that the Liouville action can be derived by counting the stress tensor gates with appropriate cost functions within Nielsen’s geometric framework.
@@
Similar optimization was also defined for the vacuum and thermal states of higher-dimensional CFTs Caputa:2017yrh, but with a prescription to optimize only over the Weyl factor of the background metric on which the CFT path integral is evaluated (unlike in 2D, this is only a subset of possible geometries in higher-dimensions). Again, it was found that optimal metrics have a constant negative Ricci scalar curvature i.e., are hyperbolic.
@@
Another idea from Caputa:2017urj, Caputa:2017yrh, dubbed path-integral complexity, was to interpret the Liouville action as a novel measure of complexity for wave functions prepared using Euclidean path integrals. The main reason to advocate this interpretation was that the on-shell Liouville action is proportional to the volume of the space on which the path integral is computed. Indeed, after fixing the normalization such that the integral of the Weyl factor $e^{2\phi}$ over a unit area (a square region of size $\epsilon$ in the original metric) has a value 1 (see detailed discussion in Bhattacharyya:2018wym), we can interpret the volume of the Euclidean space as the number of tensors in the optimal network. The on-shell action associated with the CFT vacuum state reads
$\mathcal{C}_{\Psi_{0}}=\text{Min}[S_{\mathcal{L}}[\phi]]=\frac{cL}{12\pi\epsilon},\qquad L=\int dx\,.$ (148)
@@
This result is also consistent with qualitative expectations and holographic conjecture Stanford:2014jda, which estimates the leading divergence in complexity of the CFT state by the spatial volume.
@@
Later on, researchers extended the path-integral optimization method to other settings. As we already mentioned, Caputa:2017urj, Caputa:2017yrh proposed a higher-dimensional generalization of the action (148). The the optimization of 2D CFTs deformed by relevant operators was studied in Bhattacharyya:2018wym, where a modified Liouville action was minimized over both metrics and local deformation couplings. The authors of Caputa:2020mgb investigated path integrals for inhomogeneous CFTs, that are usually interpreted as 2D CFTs on non-trivial background metrics. From the perspective of the optimization, the starting point is the CFT on a non-trivial background metric ($\hat{g}_{ab}$ instead of $\delta_{ab}$) and then consider $g_{ab}=e^{2\phi}\hat{g}_{ab}$ as the metric for optimization. This leads to additional background Ricci curvature term $\hat{R}$ in the Liouville equation which can nevertheless be solved using standard uniformization techniques.
@@
Soon after Caputa:2017urj, reference Czech:2017ryf argued that a variation of path integral complexity in holographic 2D CFTs leads to Einstein’s equations in the bulk of AdS3. Moreover, the Liouville action (or its covariant version, called the Polyakov action) motivated the treatment of Nielsen’s complexity by introducing trajectories on the unitary manifold generated by sequences of 2D conformal transformations Caputa:2018kdj (see Sec. 5.6.2).
@@
The formulation in terms of a Liouville action also inspired a series of developments in the tensor-network community. They used the path integral approach to cMERA to define tensor networks in terms of the background geometry on which the state is prepared Milsted:2018yur, Milsted:2018san. In particular, they showed how the background metric parameters translate into details of the tensor network. This idea was later used in Camargo:2019isp to argue for a Nielsen-type interpretation of the Liouville action in terms of an appropriate choice of the cost function on the complexity geometry. As opposed to the developments on Nielsen’s complexity discussed in Sec. 5, the work Camargo:2019isp necessarily utilized circuits that contained both unitary and Hermitian gates associated with components of the energy-momentum tensor operator. This is similar in spirit to the results discussed in Sec. 7.5.1 of Heller:2024ldz and has the same origin: Euclidean time evolution is an efficient and natural method for preparing useful states. This points towards a mathematical generalization of the Nielsen’s complexity that allows for exponentiation of anti-Hermitian (giving rise to unitary transformations), as well as Hermitian matrices (giving rise in particular to Euclidean time evolution).
@@
Finally, the results in Ref. Camargo:2019isp pointed out that a general cost function for path integral optimization will also contain positive powers of the cut-off $\epsilon$ associated with higher derivatives of $\phi$ in the spirit of an effective field theory. However, the solutions of Eq. (147) acquire derivatives that are themselves of the order of $1/\epsilon$, breaking the hierarchy within these more general cost functions that left the contribution (146) as the dominant term. Thus we need a more systematic inclusion of the UV cut-off $\epsilon$, bringing us to the subject of the next section.